Suppose that you estimate a multiple regression model using OLS using a sample of 60 observations. The skewness of the residuals is 0.5 and the excess kurtosis is 1. As a result, the value of the Jarque-Bera test is _______ and we ______ reject the null hypothesis of normally distributed disturbances at the 5-percent level of significance.

10; cannot

10; can

5; cannot

5; cannot
Answer #1

Answer - 5; cannot

2 ANSWER =) JB Static p-values ♡ ( skewness? + Excess Kustosis 4 6 N=60, Skewness=o-5, Eucess Kustosis = 1 p-value= 60 (co.s²